A 4 kg mass is connected to a 7 kg mass over a massless frictionless pulley as shown (Atwood’s machine) Find the acceleration of the 4 kg mass.
1
Expert's answer
2020-11-27T00:40:37-0500
"(m+M)a=(M-m)g\\(4+7)a=(7-4)9.8\\a=2.7\frac{m}{s^2}"